# Open-vocabulary Temporal Action Localization using VLMs

### v3

@@ Abstract
Video action localization aims to find timings of a specific action from a long video. Although existing learning-based approaches have been successful, those require annotating videos that come with a considerable labor cost. This paper proposes a learning-free, open-vocabulary approach based on emerging off-the-shelf vision-language models (VLM). The challenge stems from the fact that VLMs are neither designed to process long videos nor tailored for finding actions. We overcome these problems by extending an iterative visual prompting technique. Specifically, we sample video frames into a concatenated image with frame index labels, making a VLM guess a frame that is considered to be closest to the start/end of the action. Iterating this process by narrowing a sampling time window results in finding a specific frame of start and end of an action. We demonstrate that this sampling technique yields reasonable results, illustrating a practical extension of VLMs for understanding videos. A sample code is available at https://microsoft.github.io/VLM-Video-Action-Localization/
@@ I Introduction
Video action localization aims to find timings of a specific action from a long video. Extracting precise timings of an action is a fundamental technique in computer-vision and other research domains, including video annotation, video editing, and automatic video collection from unlabelled datasets. For example, in the context of robotics, accurately extracting videos corresponding to specific actions will greatly benefit from analyzing human action in several robot-teaching frameworks where robot learns from human demonstrations, such as Learning from Demonstration (LfD)  [1] and Learning from Observation (LfO) [2, 3, 4].
Previous mainstream research on video action localization has primarily focused on training models based on labelled datasets. The typical approach is to classify frames or a series of consecutive frames with predefined action labels for determining which action is being executed at each point in time in the video [5, 6, 7]. While these methods have been successful on several benchmarks, they have limitations in terms of extensibility. Specifically, if new actions not included in the training data need to be recognized, additional data collection and model retraining are required. This presents a bottleneck for the quick and convenient application of these methods to data outside benchmark tasks.
In this paper, we introduce a pipeline for Open-vocabulary video action localization (OVAL), which does not require pre-training and can address action labels in an open-vocabulary form (Fig. 1). Our approach is inspired by a recently proposed method called “Prompting with Iterative Visual Optimization (PIVOT)”[8]. The original PIVOT aims to utilize Vision-Language Models (VLM) to reason an embodied action by framing the problem as iterative visual question answering. In this paper, we propose a simple yet effective approach to apply PIVOT to find a temporal segment of actions. Specifically, we sample video frames into a concatenated image with frame index labels, making a VLM guess the frame closest to the start and end of an action. Iterating this process by narrowing the sampling time window results in finding specific frames when the action started and ended, thus achieving action localization. Although the recent advancements in VLM technology has supported the development of OVAL [9, 10, 11, 12, 13, 14], to the best of our knowledge, there are no existing methods that utilize off-the-shelf VLMs directly.
@@
Using OpenAI’s GPT-4o [15] as an example of VLMs, we evaluated the performance of the proposed pipeline using an existing benchmark of breakfast cooking dataset and our labelled original cooking dataset. We checked how the pipeline works for a typical robot-teaching condition in which a transition of high-level task needs to be grounded against a human demonstration video [16, 17, 18] (e.g., grasp, pick-up, move, put, and release an object). As result, we show the model achieves over 60% accuracy in terms of mean-over-frame (MoF), performing better than a baseline method. Although the performance appeared to be inferior to the state-the-of-art learning-based methods, the result demonstrated the feasibility of the proposed method. Additionally, we show the impact of several parameters of visual prompting and qualitative results for extracting specific actions from an unlabeled dataset, to deepen insights into its application. The contributions of this research are:
  • Proposing a novel pipeline for open-vocabulary video action localization by extending PIVOT.
  • Demonstrating its effectiveness using existing benchmarks and our own benchmarks for robot teaching
@@
The primary advantage of this method lies in its use of open vocabulary, enabling action localization across diverse contexts. We believe that its adaptability extends beyond robot teaching to various research fields and practical applications.
@@ II-A Model-based Action Localization
Model-based temporal action localization can be broadly categorized into anchor-based methods and boundary-based methods. Anchor-based methods, also known as top-down methods, mainly use multi-scale temporal sliding windows to detect pre-trained actions within untrimmed videos [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. In contrast, boundary-based methods, also known as bottom-up methods, initially detect the temporal boundaries of actions, followed by further analysis of the proposal regions estimated from these boundaries [31, 6, 7, 27, 32, 33]. These methods typically involve a two-stage pipeline consisting of temporal proposal generation and action classification [25, 27, 28, 26], while end-to-end methods have also been proposed [34, 19, 35, 36]. The proposed method follows a boundary-based approach, as it aims to detect the start and end boundaries of actions. Additionally, unlike other methods, our VLM-based method accommodates open-vocabulary queries. This allows for the addition of new actions without the need for data re-collection or re-training.
@@ III-B Implementation of T-PIVOT
The Algorithm 1 shows the specific pipeline we implemented for the experiment sin this paper. The computation yields the start and end times of a set of tasks in a video by applying T-PIVOT in parallel. We used OpenAI’s GPT-4o [15] for all the experiments. The sampling window is halved with each iteration, and this process is repeated for a specific number of iterations, which we set to four as a practical number.
@@
  • $SampleFrames$: Creates a grid of frames around $center\_time$ with an interval of $interval$ and returns the frames and their corresponding time.
  • $QueryVLM$: Uses a VLM to analyze the image and determine a frame closest to the completion of the task.
@@ IV Experiment
We evaluated the performance of the proposed pipeline in the light of practical use cases. Specifically, the pipeline was evaluated in the context of a typical robot-teaching condition on top of task and motion planning (TAMP)[39], where a system needs to ground high-level task plan (e.g., grasp, pick-up, move, put, and release an object) against a human demonstration video [16, 17, 18]. To this end, we used an existing benchmark dataset and our labeled original cooking dataset. We focused on cooking actions because it reflects realistic scenarios as well as challenging appearance of cluttered objects in a scene.
@@ IV-A Benchmark
Breakfast Dataset [40] is the third-person videos containing ten activities related to breakfast preparation, such as pouring drink and making a sandwich. Each video is annotated without gaps along the timeline with more detailed action labels, such as taking a cup and stirring. The videos contain 52 different individuals across 18 different kitchens, totaling 1712 videos. As we are proposing learning-free approach, we used only the testing dataset of 335 videos based on the split used in [41].
Fine-grained Breakfast dataset is the first-person videos of human cooking activities with fine-grained robotics tasks. Existing datasets of human actions including the Breakfast dataset [40] are labeled with coarse categories and simple motion labels often consist of multiple robotics manipulations (e.g., a motion labeled as “pour milk” needs a sequential robotic manipulation of moving a robot hand towards a milk carton, grasp it, pick it up, move it near a cup, rotate and hold the milk carton, adjust its pose, move back to the table and release it [42, 43]. There is a huge gap between the granularity of action in the computer-vision community and the robotics community. To our knowledge, however, no existing datasets label tasks based on fine-grained robotic actions. To gain insights into the usefulness of the proposed pipeline for robot teaching, we prepared a fine-grained breakfast dataset by manually annotating an existing dataset of breakfast preparation [44] using a third-party annotation tool [45].
Based on the Kuhn-Tucker theory [46] explaining solution space of linear simultaneous inequalities in space, we defined tasks as a transition of constraints imposed to an object [47, 48, 49]. As such we obtain a task set that is sufficient and enough to explain manipulation, such as picking up, moving, placing, rotating (refer to  [48] for the complete list of tasks). 113 videos were used in this study.
@@ IV-B Task Definition
Given a video $V$ and a sequence of task labels $L_{i}$ ($i=1\ldots N$) contained within it, our goal is to estimate the timing of these transitions $T_{i\rightarrow i+1}$ ($i=1\ldots N-1$). We assume that the task sequence is continuous. As a baseline method, we partition the video uniformly based on the length of the sequence, allocating equal time intervals to each task. We evaluate our method using the following metrics commonly used in temporal action localization. [36]:
@@
  • F1-score: The harmonic mean of precision and recall, providing a balanced measure of prediction accuracy.
@@ IV-C T-PIVOT for Recognizing Seamless Action Transitions
By applying the T-PIVOT method described in Section III, we estimated the timing of task transitions within the video as follows. First, the start time of each task, $T_{i\_start}$ ($i=1\ldots N$), is estimated using the T-PIVOT in parallel processing. The end time of each task, $T_{i\_end}$ ($i=1\ldots N$), is similarly estimated. This approach yields an estimated series of start times and end times for each task. In this experiment, we assume that tasks are continuously present in the video. Therefore, the end time of one task is identical to the start time of the subsequent task. To ensure this, we calculate this time as the midpoint between the estimated end time of the current task and the estimated start time of the next task:
@@ IV-D Action-order Aware Prompting
where {task_sequence} lists the sequence of tasks in the format “1. $L_{1}$, 2. $L_{2}$, …” and {task_focus} refers to the specific task being searched for in this instance (i.e., “k. $L_{k}$”). This prompt is designed to consider the order and focus of tasks during the process, making the VLM aware of which task is of the current interest.w
@@ IV-E Performance Across Different Prompting Styles
First, we investigated how to create images for inclusion in the prompt, specifically focusing on how many images from a time window should be presented and in what manner. Table I and Table II show the impact of different numbers of images, depending on the grid size, on the results. The highest performance was achieved with a 5x5 grid (i.e., extracting 25 frames from a sampling time window) across the three metrics for the Breakfast Dataset and 3x3- or 4x4-grid settings for the Fine-grained Breakfast Dataset. These results appear to highlight the trade-off between temporal and spatial resolution in this method—a smaller grid size appears to decrease the temporal resolution in recognition as it limits the number of frames that can be extracted from the video, while a larger grid size can affect the spatial resolution of each frame as the resolution of the prompt image is fixed.
We also investigated other parameters for visual prompting, including the rendering position of the annotation numbers and the tiling spacing (Fig. 4). However, these variations did not lead to noticeable performance improvement (Table III); the inferior performance observed with center-rendering of the annotation might be due to overlapping annotated numbers reducing the image clarity. Furthermore, we explored conditions where the T-PIVOT iterations were increased until consecutive frames in the video were sampled, considering the possibility that the fixed number of iterations may limit the temporal resolution. However, this approach also did not lead to notable improvement in the results (data not shown), while taking longer time to complete the inference.
TABLE I: Performance Metrics Across Grid Sizes (Breakfast Dataset)
@@
TABLE II: Performance Metrics Across Grid Sizes
@@
TABLE III: Performance Metrics Across Rendering Styles (Breakfast Dataset)
Method Original Center annotation Spacing
MoF 60.1 54.1 58.0
IoU 40.8 37.8 40.0
F1 51.2 48.1 50.5
@@ IV-F Comparison with Existing Methods
Table IV compares our proposed method with an existing method [36] on the Breakfast Dataset. It is crucial to emphasize that this comparison is not direct due to fundamental differences in experimental assumptions. Our method presupposes knowledge of the sequence of actions in the video, whereas general action localization lacks this prior information, requiring independent action labeling for each frame. The absence of predefined action sequences makes general action localization more challenging than our setup. Despite these methodological disparities, existing approaches still surpass our proposed method, underscoring its limitations compared to state-of-the-art learning-based techniques.
TABLE IV: Comparison with existing action localization
Breakfast Dataset
MoF IoU F1
Bueno-Benito et al. [36] 65.1 52.1 54.6
Temporal PIVOT 59.8 40.3 50.6

### v4

# Open-Vocabulary Action Localization with Iterative Visual Prompting

@@ Abstract
Video action localization aims to find the timings of specific actions from a long video. Although existing learning-based approaches have been successful, they require annotating videos, which comes with a considerable labor cost. This paper proposes a learning-free, open-vocabulary approach based on emerging off-the-shelf vision-language models (VLMs). The challenge stems from the fact that VLMs are neither designed to process long videos nor tailored for finding actions. We overcome these problems by extending an iterative visual prompting technique. Specifically, we sample video frames and create a concatenated image with frame index labels, making a VLM guess a frame that is considered to be closest to the start and end of the action. Iterating this process by narrowing a sampling time window results in finding the specific frames corresponding to the start and end of an action. We demonstrate that this technique yields reasonable performance, achieving results comparable to state-of-the-art zero-shot action localization. These results illustrate a practical extension of VLMs for understanding videos. Sample code is available at https://microsoft.github.io/VLM-Video-Action-Localization/
@@ I Introduction
Video action localization aims to find the timings of specific actions from a long video. Extracting precise timings of an action is a fundamental technique in computer vision and related research domains, including video annotation, video editing, and automatic video collection from unlabeled datasets. For example, in the context of robotics, accurately extracting videos corresponding to specific actions will greatly benefit from analyzing human action in several robot-teaching frameworks where robots learn from human demonstrations, such as Learning from Demonstration (LfD)  [1] and Learning from Observation (LfO) [2, 3, 4].
Previous mainstream research on video action localization has primarily focused on training models based on labeled datasets. The typical approach is a frame-by-frame method where a pipeline classifies frames or a series of consecutive frames with predefined action labels  [5, 6, 7]. While these methods have been successful on several benchmarks, they have limitations in label extensibility. Specifically, these methods require additional data collection and model retraining when new labels are added, which is a bottleneck for the quick and convenient application of these methods to data outside benchmark tasks.
In this paper, we introduce a pipeline for Open-vocabulary Video Action Localization (OVAL), which does not require pre-training and can address action labels in an open-vocabulary form (Fig. 1). Our approach is inspired by a recently proposed method called Prompting with Iterative Visual Optimization (PIVOT) [8]. The original PIVOT aims to utilize Vision-Language Models (VLMs) to reason an embodied action by framing the problem as iterative visual question answering. In this paper, we propose a simple yet effective approach to apply PIVOT to find temporal segments of actions. Specifically, we sample video frames and create a concatenated image with frame index labels, making a VLM guess the frame closest to the start and end of an action. Iterating this process by narrowing the sampling time window results in finding specific frames when the action started and ended, thus achieving action localization. Although the recent advancements in VLM technology have supported the development of OVAL [9, 10, 11, 12, 13, 14], these methods still rely on domain-specific pre-training and to the best of our knowledge, there are no training-free methods that utilize off-the-shelf VLMs directly.
@@
Using OpenAI’s GPT-4o [15] as an example of VLMs, we evaluated the performance of the proposed pipeline on an existing breakfast-cooking dataset and our own labeled cooking dataset. We checked how the pipeline works for a typical robot-teaching condition in which transitions between high-level tasks need to be grounded to a human demonstration video [16, 17, 18] (e.g., grasp, pick up, move, put, and release an object). As result, we show that the model achieved over 60% accuracy in terms of mean-over-frame (MoF), outperforming a baseline method. Although the performance appeared to be inferior to existing learning-based methods, the results demonstrated the feasibility of the proposed method. Furthermore, we tested the proposed pipeline for a general action-localization task using a sport video benchmark and showed that it performed comparably to state-of-the-art zero-shot methods. In addition, to offer further insights into its application, we explored the impact of several visual prompting parameters and present qualitative results for extracting specific actions from an unlabeled dataset. The contributions of this research are:
  • Proposing a novel pipeline for OVAL by extending PIVOT.
  • Demonstrating its effectiveness using existing benchmarks and our own benchmarks for robot teaching.
@@
The primary advantage of this method is its training-free nature for OVAL, which enables action localization across diverse contexts. We believe its adaptability extends beyond robot teaching, making it applicable to a wide range of research fields and practical use cases.
@@ II-A Model-based Action Localization
Model-based temporal action localization can be broadly categorized into anchor-based methods and boundary-based methods. Anchor-based methods, also known as top-down methods, mainly use multi-scale temporal sliding windows to detect pre-trained actions within untrimmed videos [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. In contrast, boundary-based methods, also known as bottom-up methods, initially detect the temporal boundaries of actions, followed by further analysis of the proposal regions estimated from these boundaries [31, 6, 7, 27, 32, 33]. Boundary-based methods methods typically involve a two-stage pipeline consisting of temporal proposal generation and action classification [25, 27, 28, 26], while end-to-end approaches have also been proposed [34, 19, 35, 36]. The proposed method follows a boundary-based approach, as it aims to detect the start and end boundaries of actions. Additionally, unlike other methods, our VLM-based method accommodates open-vocabulary queries. This allows for the addition of new actions without the need for data collection or re-training.
@@ III-B Implementation of T-PIVOT
The Algorithm 1 shows the specific pipeline we implemented for the experiment in this paper. The computation yields the start and end times of a set of tasks in a video by applying T-PIVOT in parallel. We used OpenAI’s GPT-4o [15] for all the experiments. The sampling window is halved with each iteration, and this process is repeated for a specific number of iterations.
@@
  • $SampleFrames$ creates a grid of frames around $center\_time$ with an interval of $interval$ and returns the frames and their corresponding time.
  • $QueryVLM$ uses a VLM to analyze the image and determine a frame closest to the completion of the task.
@@ IV Experiment
We evaluated the performance of the proposed pipeline in the light of practical use cases. Specifically, the pipeline was evaluated in the context of a typical robot-teaching condition on top of task and motion planning (TAMP)[39], where a system needs to ground high-level task plans (e.g., grasp, pick up, move, put, and release an object) to a human demonstration video [16, 17, 18]. To this end, we used an existing benchmark dataset and our own labeled cooking dataset. We focused on cooking actions because it reflects realistic scenarios as well as challenging appearance of cluttered objects in a scene.
@@ IV-A Benchmark
Breakfast Dataset [40] consists of third-person videos containing ten activities related to breakfast preparation, such as pouring a drink and making a sandwich. Each video is annotated without gaps along the timeline with more detailed action labels, such as taking a cup and stirring. The videos contain 52 different individuals across 18 different kitchens, totaling 1712 videos. As we are proposing a learning-free approach, we used only the testing dataset of 335 videos based on the split used in [41].
Fine-grained Breakfast dataset consists of the first-person videos of human cooking activities with fine-grained robotics tasks. Existing datasets of human actions including the Breakfast dataset [40] are labeled with coarse categories and simple motion labels often consist of multiple robotics manipulations. For example, a motion labeled as “pour milk” needs a sequential robotic manipulation of moving a robot hand towards a milk carton, grasp it, pick it up, move it near a cup, rotate and hold the milk carton, adjust its pose, move back to the table and release it [42, 43]—there is a huge gap between the granularity of action in the computer-vision community and the robotics community. To our knowledge, however, no existing datasets label tasks based on fine-grained robotic actions. To gain insights into the usefulness of the proposed pipeline for robot teaching, we prepared a fine-grained breakfast dataset by manually annotating an existing dataset of breakfast preparation [44] using a third-party annotation tool [45].
Based on the Kuhn-Tucker theory [46] explaining solution space of linear simultaneous inequalities in space, we defined tasks as transitions of constraints imposed on an object [47, 48, 49]. As such we obtain a task set that is sufficient and exhaustive to explain manipulation, such as picking up, moving, placing, rotating (refer to  [48] for the complete list of tasks). 113 videos were used in this study.
@@ IV-B Task Definition
Given a video $V$ and a sequence of task labels $L_{i}$ ($i=1\ldots N$) contained within it, our goal is to estimate the timing of these transitions $T_{i\rightarrow i+1}$ ($i=1\ldots N-1$). We assume that the task sequence is continuous. As a baseline method, we partition the video uniformly based on the length of the sequence, allocating equal time intervals to each task. We evaluate our method using the following metrics commonly used in temporal action localization [36]:
@@
  • F1-score: represents the harmonic mean of precision and recall, providing a balanced measure of prediction accuracy.
@@ IV-C T-PIVOT for Recognizing Seamless Action Transitions
By applying the T-PIVOT method described in Section III, we estimated the timing of task transitions within the video as follows. First, the start time of each task, $T_{i\_start}$ ($i=1\ldots N$), is estimated using the T-PIVOT in parallel processing. The end time of each task, $T_{i\_end}$ ($i=1\ldots N$), is similarly estimated. This approach yields an estimated series of start and end times for each task. In this experiment, we assume that tasks are continuously present in the video. Therefore, the end time of one task is identical to the start time of the subsequent task. To ensure this, we calculate this time as the midpoint between the estimated end time of the current task and the estimated start time of the next task:
@@ IV-D Action-order Aware Prompting
where {task_sequence} lists the sequence of tasks in the format “1. $L_{1}$, 2. $L_{2}$, …” and {task_focus} refers to the specific task being searched for in this instance (i.e., “k. $L_{k}$”). This prompt is designed to consider the order and focus of tasks during the process, making the VLM aware of which task is of the current interest.
@@ IV-E Performance Across Different Prompting Styles
We first investigated the effectiveness of the iterative process on performance. Table I shows the impact on performance of varying the number of iterations, using the Breakfast Dataset. In this experiment, we used a 2x2 grid configuration, which led to coarse temporal resolution per iteration. The results showed that overall performance improved over the iterations, suggesting that the iterative process effectively enhances sampling resolution. Notably, even with the coarse 2x2 grid setting, performance peaked within four iterations. For subsequent experiments, the number of iterations was set to four as a practical choice.
Second, we investigated how to create images for inclusion in the prompt, specifically focusing on how many images from a time window should be presented and in what manner. Table II and Table III show the impact of grid size, which reflects the number of sampled images per iteration, on the results. The highest performance was achieved with a 5x5 grid (i.e., extracting 25 frames from a sampling time window) across the three metrics for the Breakfast Dataset, and 3x3- or 4x4-grid settings for the Fine-grained Breakfast Dataset. These results appear to highlight the trade-off between temporal and spatial resolution in this method—–a smaller grid size reduces the temporal resolution in recognition as it limits the number of frames that can be extracted from the video, while a larger grid size lowers the spatial resolution of each frame, assuming the number of iterations and the resolution of the prompt image remain the same.
Lastly, we investigated other parameters for visual prompting, including the rendering position of the annotation numbers and the tiling spacing (Fig. 4). We also explored an approach where the images were input into the model as stacked buffers rather than tiled. The highest performance across the three metrics was achieved using either the original tiling method or the stacking method. The inferior performance observed with center-rendering of the annotation might be due to overlapping annotated numbers reducing the image clarity.
TABLE I: Performance Metrics Across Iteration Counts (Breakfast Dataset)
Iteration count 1 2 3 4 5
MoF 32.4 39.7 39.5 40.7 38.6
IoU 18.4 23.8 24.0 24.5 23.2
F1 24.4 31.6 31.9 32.3 30.7
Images were rendered with 2x2 grid tiling.
TABLE II: Performance Metrics Across Grid Sizes (Breakfast Dataset)
@@
TABLE III: Performance Metrics Across Grid Sizes
@@
TABLE IV: Performance Metrics Across Rendering Styles (Breakfast Dataset)
Method Original Center annotation Spacing Stacking
MoF 60.1 54.1 58.0 59.6
IoU 40.8 37.8 40.0 45.4
F1 51.2 48.1 50.5 54.8
@@ IV-F Comparison with Existing Methods
Table V compares our proposed method with existing methods on the Breakfast Dataset, with the reported numbers obtained from [36]. It is important to note that this comparison is not direct due to fundamental differences in experimental assumptions. T-PIVOT assumes knowledge of the action sequence in the video for task-transition estimation, while learning-based methods typically do not rely on this information, making them more challenging. Despite these differences, existing methods still outperform our approach, highlighting its limitations compared to state-of-the-art learning-based techniques.
To gain insights into the zero-shot capabilities of our method, we compared its performance with existing zero-shot action localization methods. To this end, we used the THUMOS14 dataset and its officially provided evaluation code [50]. Importantly, this task is not transition estimation because the THUMOS14 dataset does not assume that tasks are continuously present in the video. Thus, we employed a different approach from the one in subsection IV-C. We divided each video into fixed time intervals and applied Algorithm 1 within those intervals to localize the start and end of actions. Specifically, the video was divided into intervals of five seconds, and a 5x5 grid tiling with four iterations was applied. To operate T-PIVOT, we assumed that the action labels contained in the video were known. Following the methods in [14], we report the mean average precision (mAP), which represents the mean of the average precision values computed at different IoU thresholds ([0.3, 0.4, 0.5, 0.6, 0.7]). Table VI shows the comparison results, where the reported numbers were obtained from [14]. These results show that the proposed method achieves state-of-the-art performance. Importantly, the existing method employs pre-training even though the performance is tested under zero-shot conditions, while our method does not require any pre-training.
TABLE V: Comparison with existing action localization methods (Breakfast Dataset)
Methods MoF IoU F1
LSTM+AL [51] 42.9 46.9 -
VTE [52] 52.2 - -
DGE [53] 59.5 48.5 51.7
TW-FINCH [54] 62.7 42.3 49.8
ABD [55] 64.0 - 52.3
Triplet [36] 65.1 52.1 54.6
T-PIVOT 60.1 40.8 51.2
The scores for the top and the second best performances are bolded and underlined, respectively.
TABLE VI: Comparison with existing zero-shot methods (THUMOS14 dataset)
Methods 0.3 0.4 0.5 0.6 0.7 Avg mAP
TMaxer [56] 10.6 9.4 8.0 6.2 4.4 7.7
ActionFormer [57] 11.3 10.0 8.4 6.6 4.6 8.2
TriDet [58] 15.2 13.2 10.8 7.9 5.2 10.5
B-II [59] 21.0 16.4 11.2 6.3 3.2 11.6
B-I [59] 27.2 21.3 15.3 9.7 4.8 15.7
Eff-Prompt [37] 37.2 29.6 21.6 14.0 7.2 21.9
STALE [60] 38.3 30.7 21.2 13.8 7.0 22.2
DeTAL [14] 38.3 32.3 24.4 16.3 9.0 24.1
OVTAL [13] - - - - - 24.9
T-PIVOT 45.4 33.6 23.2 15.5 11.4 25.8
The scores for the top and the second best performances are bolded and underlined, respectively.